301 views
3 3 votes

Three fair coins are flipped. Given that at least one of the coins shows a head and at least one of the coins shows a tail, what is the probability that exactly two of the coins show heads?

  1. $\dfrac{1}{4}$
     
  2. $\dfrac{1}{3}$
     
  3. $\dfrac{1}{2}$
     
  4. $\dfrac{2}{3}$

4 Answers

3 3 votes
When $3$ fair coins are flipped, the total sample space is

$S=\{HHH,\ HHT,\ HTH,\ THH,\ HTT,\ THT,\ TTH,\ TTT\}.$

We are given that $\textit{at least one coin shows a head and at least one coin shows a tail}.$

So the outcomes $HHH$ and $TTT$ are not possible.

Hence, the reduced sample space is

$S'=\{HHT,\ HTH,\ THH,\ HTT,\ THT,\ TTH\}.$

Thus, the number of possible outcomes is

$n(S')=6.$

Now we need the probability that $\textit{exactly two of the coins show heads}.$

The outcomes with exactly two heads are

$\{HHT,\ HTH,\ THH\}.$

So the number of favorable outcomes is $3.$

$\therefore P(\text{exactly two heads} \mid \text{at least one head and at least one tail})=\frac{3}{6}=\frac{1}{2}.$

Answer: $\boxed{\dfrac{1}{2}}$
1 1 vote

Given, 

3 fair coins are flipped with at least 1H and 1T.

So, Sample Space S = {HHT, HTH, HTT, THH, THT, TTH} (total of 6 outcomes) since HHH and TTT violates the condition of atleast 1H and 1T.

Now, 

Conditions satisfying atleast 2H are: HHT, HTH, THH (3 outcomes).

So, 3/6 = 1/2

Option C is the correct answer.

0 0 votes

Let A be the event of exactly two coins showing heads and 

B be the event of  atleast one coin showing head and atleast one coin showing tail
 

Now, we have to find out $\textbf{P(A | B)}$

$P(A | B )  = \frac{P(A,B)}{P(B)}$

$P(A,B) = \frac{3}{2^3} = $ P(2 heads, 1 tail)

$P(B) = $ P(atleast one head, atleast one tail) = 1 - P(0 heads) - P(0 tails) = $1 - \frac{1}{8} - \frac{1}{8} = \frac{6}{8}$ 

Now, $P(A | B ) = \frac{\frac{3}{2^3}}{\frac{6}{8}} = \frac{3}{6} = \frac{1}{2}$

Hence, Answer is Option C

Answer:
Position:
Show:

Related questions

5 5 votes
2 2 answers
219
219 views
GO Classes asked Apr 9
219 views
A medical laboratory buys testing kits from two manufacturers, $A$ and $B$. Manufacturer $A$ supplies $65\%$ of the kits, and manufacturer $B$ supplies $35\%$ of the kits...
3 3 votes
2 2 answers
264
264 views
GO Classes asked Apr 9
264 views
Let $S$ be a sample space and let $A$, $B$, and $C$ be three mutually exclusive events such that $A \cup B \cup C = S.$If $P(A)=x$, $P(B)=2x$, and $P(C)=1-3x$, then the m...
3 3 votes
3 3 answers
254
254 views
GO Classes asked Apr 9
254 views
A deck of $6$ cards, each carrying a distinct number from $1$ to $6$, is shuffled thoroughly. Two cards are then removed one at a time from the deck. What is the probabil...
4 4 votes
2 2 answers
275
275 views
GO Classes asked Apr 9
275 views
The probability that a student passes Algebra is $0.7$. The probability that the student passes Biology is $0.6$. The probability that the student passes Chemistry is $0....